# SCREP: Scene Coordinate Regression and Evidential Learning-based Perception-Aware Trajectory Generation

### v1

@@ Abstract
Autonomous flight in GPS‑denied indoor spaces requires trajectories that keep visual‑localization error tightly bounded across varied missions. Whereas visual inertial odometry (VIO) accumulates drift over time, scene coordinate regression (SCR) yields drift‑free, high‑accuracy absolute pose estimation. We present a perception‑aware framework that couples an evidential learning‑based SCR pose estimator with a receding horizon trajectory optimizer. The optimizer steers the onboard camera toward pixels whose uncertainty predicts reliable scene coordinates, while a fixed-lag smoother fuses the low-rate SCR stream with high-rate IMU data to close the perception–control loop in real time. In simulation, our planner reduces translation (rotation) mean error by 54% / 15% (40% / 31%) relative to yaw-fixed and forward-looking baselines, respectively. Moreover, hardware‑in‑the‑loop experiment validates the feasibility of our proposed framework.
@@ I Introduction
Vision-based localization has made rapid progress in recent years, allowing unmanned aerial vehicles (UAVs) to fly autonomously in GPS-denied indoor environments. Most perception-aware trajectory planning approaches, [1, 2, 3], steer the camera toward feature-rich regions to improve the pose estimate obtained by visual inertial odometry (VIO). Although effective at reducing drift over short distances, these methods are still limited by the cumulative error of VIO.
An alternative is to employ scene coordinate regression (SCR), a learning-based method predicting a 3D coordinate for every pixel [4, 5, 6, 7, 8] instead of VIO. A Perspective‑n-Point (PnP) solver then produces a drift‑free absolute camera pose, achieving high metric accuracy even over long trajectories. However, not all pixels produce reliable coordinates. Moreover, SCR poses have both outliers and a low frequency latency, which complicates their use with real-time planning and control.
This paper proposes a perception-aware trajectory planning framework that explicitly utilizes the pixel-wise uncertainty of an evidential learning-based SCR network (E-SCRNet). To the best of our knowledge, this is the first active-perception research using SCR. We train a neural network for SCR with evidential learning [9], so each pixel outputs scene coordinates and both aleatoric and epistemic uncertainties. We use the entropy as the uncertainty metric since it jointly captures aleatoric and epistemic uncertainties. The entropy map drives a receding horizon trajectory optimizer that (i) weights low-entropy pixels higher and (ii) removes high-entropy pixels from the cost, thereby orienting the camera toward informative viewpoints while following the given reference position trajectory. A fixed-lag smoother then fuses the low-rate SCR pose with high-rate IMU measurements, reducing latency, suppressing outlier SCR pose estimates and eliminating the drift caused by IMU bias.
@@
Figure 1: Top‑down 3‑D reconstructions of the two sites produced by SCR. (left) RGB-colored points (right) normalized entropy map.
Figure 2: Illustration of proposed receding horizon perception‑aware trajectory generation. (blue) Camera orientations are optimized to steer the camera toward low entropy coordinates. (red) position trajectory (black) current pose.
The main contributions of this research are as follows:
  • We present a real‑time and receding-horizon framework that injects E-SCRNet uncertainty into perception-aware trajectory optimization, sustaining high-accuracy localization.
@@
II Related Work
II-A Perception–Aware Trajectory Planning
PAMPC [1] augments nonlinear model predictive control (MPC) with a visibility cost and Zhang et al. [2] frame perception–aware planning in a receding-horizon fashion. Yaw optimization was later coupled with differential flatness and a differentiable camera cost [3], extended to incorporate semantic priors [10], time‑optimal parameterization [11]. Other directions include Fisher information [12], topological path planning [13]. Recent studies combine position–yaw optimization with a novel feature‑visibility model [14], feature‑limited or unknown environments [15] and introduces control barrier functions for safety guarantees [16]. However, all of the above rely on visual–inertial odometry (VIO), and the accumulated drift of VIO is merely reduced rather than eliminated.
II-B Learning-based Visual Localization
Beyond such model-based visual localization methods, numerous learning-based visual localization methods have emerged. Absolute Pose Regressors (APRs) employ neural networks to directly regress camera pose [17, 18, 19]. Uncertainty‑aware variants [20, 21] also estimate uncertainty, yet APRs remain markedly less accurate than both model‑based pipelines and other learning‑based alternatives [22].
HLoc [23] retrieves candidate keyframes with global descriptors and refines the pose via local feature matching. While accurate, its memory and computation grow with the number of keyframes, being likely to exceed the resources available on onboard hardware as the environment expands.
SCR regresses a dense 3‑D scene‑coordinate map and then estimates the pose with PnP-RANSAC. Early work used random forests [4]. DSAC replaced it with differentiable RANSAC and CNNs [5, 6]. Hierarchical SCR has recently been explored for large scenes [7, 8]. SCR attains high absolute accuracy with a single forward pass but can produce outliers.
II-C Uncertainty Quantification
Bayesian neural networks (BNNs) estimate aleatoric and epistemic uncertainty via Monte‑Carlo dropout [24], deep ensembles [25], or more general Bayesian learning frameworks [26]. Uncertainty‑aware APR [20] and VIO using BNNs and neural radiance field [27] show the benefit of uncertainty for localization. However, BNNs require multiple forward passes for the uncertainty quantification.
On the other hand, evidential learning offers closed‑form aleatoric and epistemic estimates without sampling [9, 28] and has been adopted for NeRF‑aided SCR [29], semantic segmentation [30], semantic mapping [31], and traversability learning [32].
@@ III System Overview
Given generic navigation scenarios, this study aims to improve an initial positional path by simultaneously refining the vehicle’s position and yaw trajectories.
Problem. Within each receding horizon of length $T_{plan}$, the UAV is given (i) reference position waypoints or trajectory, (ii) regressed scene–coordinates and their pixel‑wise uncertainties, and (iii) the current UAV pose estimate. The objective is to produce a dynamically feasible trajectory that (a) follows the reference, (b) respects velocity/acceleration limits, and (c) orients viewpoints with low uncertainty to maximize localization accuracy. We assume (A1) lower uncertainty implies higher scene‑coordinate accuracy and (A2) uncertainty varies smoothly, allowing its use within short horizons. Thanks to the differential flatness of multirotor dynamics [33], the position and yaw trajectories $\{\mathbf{p}(t),\psi(t)\}$, and their derivatives fully determine the required state and control.
Pipeline. Fig. 3 depicts the pipeline for the receding horizon trajectory generation achieving the objective. Images from a camera pass through E‑SCRNet [29] to produce scene coordinates and uncertainties, which a PnP‑RANSAC module converts into camera poses. To compensate for the latency and outliers of this stream, we fuse it with high‑rate IMU measurements in a fixed‑lag GTSAM smoother [34], yielding a clean, high‑frequency state estimate. Note that other sensor fusion methods, such as the Kalman Filter, would also suffice for the UAV pose estimation.
The pose estimate, waypoints, scene coordinates, and uncertainties feed a perception‑aware optimizer, split into position and yaw sub‑problems for real‑time execution. The resulting reference trajectories $\{\mathbf{p}^{\star}(t),\psi^{\star}(t)\}$ is repeatedly generated until the UAV reaches its goal pose.
@@ IV Methodology
In this section, we first describe the training of E‑SCRNet, a SCR network trained via deep evidential regression [9], and discuss its representation of uncertainty for the perception-aware trajectory planning. We then present our B‑spline–based, receding horizon, perception-aware trajectory planner utilizing these scene coordinates and their associated uncertainties.
@@ IV-A Scene Coordinate Regression with Deep Evidential Regression
In this study, we train a neural network for SCR and assess the uncertainty of each predicted scene coordinate so that the perception‑aware trajectory optimization can reason about perception quality. We adopt the deep evidential regression [9], following the approach in [29], as it can jointly quantify both aleatoric and epistemic uncertainties without requiring multiple forward passes, unlike BNNs. Note that the architecture of the SCR network is tangential to this work.
@@
Let $\mathbf{v}\in\mathbb{R}^{3}$ denote the scene coordinate of each pixel. For brevity, we write $v$ when referring to one Cartesian component. We model $v\sim\mathcal{N}(\mu,\sigma^{2})$ where the unknown mean and variance, $\mathbf{t}=(\mu,\sigma^{2})$, are themselves random variables drawn from a Gaussian and inverse gamma prior, $\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\lambda\bigr{)}$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$. Here, $\Gamma(\cdot)$ is the gamma function. Assuming that the priors on $\mu$ and $\sigma^{2}$ factorize each other, $p(\mathbf{t}|\mathbf{m})$ is the Normal Inverse-Gamma (NIG) distribution, which is a Gaussian conjugate prior, parameterized by $\mathbf{m}=(\gamma,\lambda,\alpha,\beta)$ with $\lambda>0$, $\alpha>1$, and $\beta>0$.
@@
This NIG distribution serves as a higher‑order evidential distribution from which the lower‑order Gaussian likelihood that generates scene coordinates is sampled, allowing the network to express the prediction $\mathbb{E}[\mu]=\gamma$, the aleatoric uncertainty $\mathbb{E}[\sigma^{2}]=\beta/(\alpha-1)$, and epistemic uncertainty $\text{Var}[\mu]=\beta/\big{(}\lambda(\alpha-1)\big{)}$.
@@
The deep evidential regression aims to learn the hyperparameters $\mathbf{m}_{\theta}=(\gamma_{\theta},\lambda_{\theta},\alpha_{\theta},\beta_{\theta})$. We can express the likelihood of the scene coordinate given the hyperparameters $\mathbf{m}_{\theta}$ as a closed-form $p(v|\mathbf{m}_{\theta})=\text{St}\big{(}v;\gamma_{\theta},\beta_{\theta}(1+\lambda_{\theta})/(\lambda_{\theta}\alpha_{\theta}),2\alpha_{\theta}\big{)}$ where $\text{St}\big{(}v;\eta,\tau^{2},\chi)$ is a Student-t distribution with location $\eta$, scale $\tau^{2}$, and degree of freedom $\chi$.
@@
The total loss is the mean of the per‑sample loss, $\mathcal{L}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\theta}^{i}$, where each sample‑wise loss is $\mathcal{L}_{\theta}^{i}=\mathcal{L}_{\theta,\mathrm{nll}}^{i}+\rho\,\mathcal{L}_{\theta,\mathrm{reg}}^{i}$. The loss $\mathcal{L}_{\theta,nll}^{i}$, is chosen as the negative log-likelihood (NLL) of $p(v|\mathbf{m}_{\theta})$ for maximum likelihood estimation (MLE):
@@
$\begin{split}\mathcal{L}_{\theta,nll}^{i}&=\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right)\\ &+\tfrac{1}{2}\log\!\left(\pi/\lambda_{\theta}\right)-\alpha_{\theta}\log\!\bigl{(}2\beta_{\theta}(1+\lambda_{\theta})\bigr{)}\\ &+\bigl{(}\alpha_{\theta}+\tfrac{1}{2}\bigr{)}\log\!\bigl{(}(v_{i}-\gamma_{\theta})^{2}+2\beta_{\theta}(1+\lambda_{\theta})\bigr{)}.\end{split}$ (1)
The loss $\mathcal{L}_{\theta,reg}$ is designed for regularization:
$\mathcal{L}_{\theta,reg}^{i}=(2\lambda_{\theta}+\alpha_{\theta})|v_{i}-\gamma_{\theta}|.$ (2)
@@
After training, we can express the predictive entropy, $H_{\theta}=\mathbb{E}_{v\sim p(v|\mathbf{m}_{\theta})}[-\log p(v|\mathbf{m}_{\theta})]$, admits the closed form
@@
$\begin{split}H_{\theta}&=\!(\alpha_{\theta}+\tfrac{1}{2})(\psi(\alpha_{\theta}+\tfrac{1}{2})-\psi(\alpha_{\theta}))\\ &+\tfrac{1}{2}\log\!\left(2\pi\beta_{\theta}(1+\lambda_{\theta})/\lambda_{\theta}\right)\\ &+\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right)\end{split}$ (3)
@@
where $\psi(\cdot)$ is the digamma function. We use the entropy as a representation of uncertainty instead of aleatoric and epistemic uncertainty since it can represent the aleatoric and epistemic uncertainty. The detailed justification will be given in Section V-B. We refer the reader to [9] for more information on deep evidential regression.

### v2

@@ Abstract
Autonomous flight in GPS‑denied indoor spaces requires trajectories that keep visual‑localization error tightly bounded across varied missions. Map‑based visual localization methods such as feature matching require computationally intensive map reconstruction and have feature‑storage scalability issues, especially for large environments. Scene coordinate regression (SCR) provides an efficient learning-based alternative that directly predicts 3D coordinates for every pixel, enabling absolute pose estimation with significant potential for onboard robotics applications. We present a perception‑aware trajectory planner that couples an evidential learning‑based SCR pose estimator with a receding-horizon trajectory optimizer. The optimizer steers the onboard camera toward reliable scene coordinates with low uncertainty, while a fixed-lag smoother fuses the low-rate SCR pose estimates with high-rate IMU data to provide a high-quality, high-rate pose estimate. In simulation, our planner reduces translation and rotation RMSE by at least 4.9% and 30.8% relative to baselines, respectively. Hardware‑in‑the‑loop experiments validate the feasibility of our proposed trajectory planner under close-to-real deployment conditions.
@@ I Introduction
Vision-based localization has made rapid progress in recent years, contributing to advances in unmanned aerial vehicle (UAV) navigation systems for GPS-denied indoor environments. Map-based visual localization approaches leverage prior environmental knowledge to provide drift-free pose estimates, making trajectory planning that considers localization performance crucial for reliable navigation in such environments. Most perception-aware trajectory planning approaches [13, 39, 30, 4, 34, 40, 24, 12, 11, 38, 23] steer the camera toward feature-rich regions to improve the pose estimate obtained by structure from motion (SfM) or visual-inertial odometry (VIO). However, SfM-based feature maps require computationally intensive reconstruction and have feature‑storage scalability issues, especially for large environments, making them challenging for onboard robotics applications. Moreover, even with perception-aware guidance, VIO-based methods are still limited by cumulative error.
@@
Figure 1: Top‑down 3‑D reconstructions of the two sites produced by SCR. (left) RGB-colored points (right) Normalized entropy map. The accumulated scene coordinates and their uncertainties from multiple viewpoints reveal spatially consistent uncertainty distributions across the environment.
Figure 2: Illustration of the proposed receding-horizon perception‑aware trajectory generation. (blue) Camera orientations are optimized to steer the camera toward low-entropy coordinates. (red) Position trajectory (black) Current pose.
Scene coordinate regression (SCR) offers a learning-based alternative that directly predicts a 3D coordinate for every pixel from a single image [33, 6, 7, 27, 37, 5, 36, 17]. A Perspective‑n-Point (PnP) solver then produces a drift‑free absolute camera pose, achieving high metric accuracy even over long trajectories while requiring only a compact neural network. However, not all pixels produce reliable coordinates. Moreover, SCR often produces outliers and runs at low frequency with high latency, which complicates its use in real-time planning and control.
This paper proposes a perception-aware trajectory planning algorithm that explicitly utilizes the pixel-wise uncertainty of an evidential learning-based SCR network (E-SCRNet). To the best of our knowledge, this is the first active-perception work using SCR. We train a neural network for SCR with evidential learning [1], so each pixel outputs scene coordinates and both aleatoric and epistemic uncertainties. We use the entropy as the uncertainty metric since it jointly captures aleatoric and epistemic uncertainties. The entropy map drives a receding-horizon trajectory optimizer that (i) weighs low-entropy pixels higher and (ii) removes high-entropy pixels from the cost, thereby orienting the camera toward informative viewpoints while following the given reference position trajectory. A fixed-lag smoother then fuses the low-rate SCR pose with high-rate IMU measurements, reducing latency, mitigating outlier SCR pose estimates, and eliminating drift caused by IMU bias.
The main contributions of this work are as follows:
  • We present a real‑time, receding-horizon trajectory planner that injects E-SCRNet uncertainty into perception-aware trajectory optimization, facilitating high-accuracy localization.
@@
II Related Work
II-A Perception-Aware Trajectory and Path Planning
PAMPC [13] augments nonlinear model predictive control (MPC) with a visibility cost and Zhang et al. [39] frame perception–aware planning in a receding-horizon fashion. Yaw optimization was later coupled with differential flatness and a differentiable camera cost [30], extended to incorporate semantic priors [4], and parameterized time‑optimally [34]. Other directions incorporate information-theoretic [40] or topological metrics [24].
More recent work combines position–yaw optimization with a novel feature‑visibility model [11], considers feature‑limited or unknown environments [38], introduces control barrier functions for safety guarantees [23], or is more data-driven [12]. However, all of these approaches rely on feature matching (FM) or visual–inertial odometry (VIO). Unlike other methods, our work addresses perception-aware planning under scene coordinate regression (SCR).
II-B Learning-based Visual Localization
Beyond FM and VIO, which utilize information from visual features, numerous learning-based visual localization methods have emerged. Absolute Pose Regressors (APRs) employ neural networks to directly regress camera pose [22, 20, 35]. Uncertainty‑aware variants [19, 29] also estimate uncertainty, yet APRs remain markedly less accurate than both model‑based pipelines and other learning‑based alternatives [32].
HLoc [31] retrieves candidate keyframes with global descriptors and refines the pose via local feature matching. While accurate, its memory and computation grow almost linearly with the number of keyframes and can exceed the resources available on onboard hardware as the map grows.
SCR regresses a dense 3‑D scene‑coordinate map and then estimates the pose with PnP-RANSAC. The evolution of SCR began with random forests [33], which DSAC later replaced with differentiable RANSAC and CNNs [6, 7] for end-to-end training. Without depth supervision, ACE [5] enables SCR to be trained in a few minutes. Recent advances include hierarchical SCR [27, 37] and a pretrained image encoder for co-visibility among images [36, 17] to handle large scenes, with [17] additionally improving robustness. Although SCR attains high absolute accuracy with a single forward pass, it can produce outliers. Also, its application in robotics remains underexplored.
II-C Uncertainty Quantification
Bayesian neural networks (BNNs) estimate aleatoric and epistemic uncertainty via Monte‑Carlo dropout [14], deep ensembles [26], or more general Bayesian learning frameworks [21]. Uncertainty‑aware APR [19] and VIO using BNNs and neural radiance fields [16] show the benefit of incorporating BNN uncertainty estimates for localization. However, BNNs require multiple forward passes for the uncertainty quantification.
On the other hand, evidential learning offers closed‑form aleatoric and epistemic estimates without sampling [1, 9] and has been adopted for NeRF‑aided SCR [10], semantic segmentation [2], semantic mapping [25], and traversability learning [8]. By applying evidential learning to our approach, the uncertainty of each predicted scene coordinate is quantified with only a single forward pass.
@@ III System Overview
Given generic navigation scenarios, this study aims to improve an initial positional path by simultaneously optimizing the vehicle’s position and yaw trajectories.
Problem. At each step, we aim to compute a path of length $T_{\mathrm{plan}}$ in a receding-horizon fashion based on the following inputs: (i) reference position waypoints or trajectory, (ii) predicted scene coordinates and their associated uncertainties from E-SCRNet, and (iii) the current UAV pose estimate. The objective is to produce a dynamically feasible trajectory that (a) follows the reference, (b) respects velocity/acceleration limits, and (c) orients viewpoints with low uncertainty to maximize localization accuracy. Thanks to the differential flatness of multirotor dynamics [28], the position and yaw trajectories $\{\mathbf{p}(t),\psi(t)\}$, and their derivatives fully determine the required state and control.
Pipeline. Fig. 3 depicts the pipeline for our receding-horizon trajectory generation framework. Images from a camera pass through E‑SCRNet [10] to produce scene coordinates and uncertainties, which a PnP‑RANSAC module converts into camera poses. To compensate for the latency and outliers of this stream, we fuse it with high‑rate IMU measurements using a fixed‑lag smoother, yielding a clean, high‑frequency state estimate.
This pose estimate, the most recent scene coordinates and uncertainties from the previous execution period $T_{\text{exec}}$, and reference waypoints feed a perception‑aware trajectory optimization. This optimization is split into position and yaw subproblems in order to achieve the runtime efficiency required for real‑time execution. The resulting reference trajectory $\{\mathbf{p}^{\star}(t),\psi^{\star}(t)\}$ is continuously generated until the UAV reaches its goal pose.
@@ IV Methodology
In this section, we first describe the training of E‑SCRNet, an SCR network trained via deep evidential regression [1], and discuss its representation of uncertainty in the context of perception-aware trajectory planning. We then present our B‑spline–based receding-horizon perception-aware trajectory planner that utilizes these scene coordinates and their associated uncertainties.
@@ IV-A Scene Coordinate Regression with Deep Evidential Regression
We train a neural network for SCR and assess the uncertainty of each predicted scene coordinate so that the perception‑aware trajectory optimization can reason about perception quality. We adopt deep evidential regression [1], following the approach in [10]. In contrast to BNNs, deep evidential regression can jointly quantify both aleatoric and epistemic uncertainty without requiring multiple forward passes. Unlike previous works [1, 10], our approach represents the total uncertainty as entropy to identify reliable scene coordinates.
@@
Let $\mathbf{v}\in\mathbb{R}^{3}$ denote the scene coordinate of each pixel. For brevity, we write $v$ when referring to one Cartesian component. We model $v\sim\mathcal{N}(\mu,\sigma^{2})$, where the unknown mean and variance $\mathbf{t}=(\mu,\sigma^{2})$ are themselves random variables drawn from a Gaussian and inverse gamma prior $\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\lambda\bigr)$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$. Assuming that the priors on $\mu$ and $\sigma^{2}$ are factorized, $p(\mathbf{t}|\mathbf{m})$ is the Normal Inverse-Gamma (NIG) distribution, which is a Gaussian conjugate prior parameterized by $\mathbf{m}=(\gamma,\lambda,\alpha,\beta)$ with $\lambda>0$, $\alpha>1$, and $\beta>0$.
@@
The NIG distribution serves as a higher‑order evidential distribution from which the lower‑order Gaussian likelihood that generates scene coordinates is sampled, allowing the network to express the prediction $\mathbb{E}[\mu]=\gamma$}, the aleatoric uncertainty $\mathbb{E}[\sigma^{2}]=\beta/(\alpha-1)$, and epistemic uncertainty $\text{Var}[\mu]=\beta/\big(\lambda(\alpha-1)\big)$.
@@
Deep evidential regression aims to learn the hyperparameters $\mathbf{m}_{\theta}=(\gamma_{\theta},\lambda_{\theta},\alpha_{\theta},\beta_{\theta})$. We can express the likelihood of the scene coordinate given the hyperparameters $\mathbf{m}_{\theta}$ in closed-form $p(v|\mathbf{m}_{\theta})=\text{St}\big(v;\gamma_{\theta},\beta_{\theta}(1+\lambda_{\theta})/(\lambda_{\theta}\alpha_{\theta}),2\alpha_{\theta}\big)$ where $\text{St}\big(v;\eta,\tau^{2},\chi)$ is a Student-t distribution with location $\eta$, scale $\tau^{2}$, and degree of freedom $\chi$.
@@
The total loss $\mathcal{L}_{\theta}$ minimized during training is the mean of the per‑sample loss $\mathcal{L}_{\theta}^{i}=\mathcal{L}_{\theta,\mathrm{nll}}^{i}+\rho\,\mathcal{L}_{\theta,\mathrm{reg}}^{i}$}. The loss $\mathcal{L}_{\theta,\mathrm{nll}}^{i}$ is chosen as the negative log-likelihood (NLL) of $p(v|\mathbf{m}_{\theta})$ for maximum likelihood estimation (MLE):
@@
$\begin{split}&\mathcal{L}_{\theta,\mathrm{nll}}^{i}=\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right)\\ &+\tfrac{1}{2}\log\!\left(\pi/\lambda_{\theta}\right)-\alpha_{\theta}\log\!\left(2\beta_{\theta}(1+\lambda_{\theta})\right)\\ &+\bigl(\alpha_{\theta}+\tfrac{1}{2}\bigr)\log\!\left(\lambda_{\theta}(v_{i}-\gamma_{\theta})^{2}+2\beta_{\theta}(1+\lambda_{\theta})\right).\end{split}$ (1)
The loss $\mathcal{L}_{\theta,\mathrm{reg}}$ is designed for regularization:
$\mathcal{L}_{\theta,\mathrm{reg}}^{i}=(2\lambda_{\theta}+\alpha_{\theta})|v_{i}-\gamma_{\theta}|.$ (2)
@@
After training, we can express the predictive entropy $H_{\theta}=\mathbb{E}_{v\sim p(v|\mathbf{m}_{\theta})}[-\log p(v|\mathbf{m}_{\theta})]$ in its closed form
@@
$\begin{split}H_{\theta}&=\!(\alpha_{\theta}+\tfrac{1}{2})(\psi(\alpha_{\theta}+\tfrac{1}{2})-\psi(\alpha_{\theta}))\\ &+\tfrac{1}{2}\log\!\left(2\pi\beta_{\theta}(1+\lambda_{\theta})/\lambda_{\theta}\right)\\ &+\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right),\end{split}$ (3)
@@
where $\psi(\cdot)$ is the digamma function. We use entropy as the uncertainty representation because it captures both aleatoric and epistemic uncertainty. The detailed justification is given in Section V-B. We refer the reader to [1] for more information on deep evidential regression.